OpenStudy (anonymous):

please help Find the mean and standard deviation of the following probability distribution: x 1 2 3 P(x) 0.2 0.4 0.4

OpenStudy (tkhunny):

E[x] = 1*0.2 + 2*0.4 + 3*0.4 = Mean = First Moment E[x^2] = 1*0.2 + 4*0.4 + 9*0.4 = Second Moment Var(X) = ??? You tell me. StdDev(X) = ???

OpenStudy (anonymous):

thank you